write your explanation of each step as you would explain the following algebraic problem to a learner 2.3^x=162
1
Expert's answer
2020-09-09T16:18:36-0400

"2\u00d73^x=162"

Divide each side by 2

"3^x=81"

We find the prime factors of 81.

"81=3^4"

Then

"3^x=3^4"

Since both sides the equation has the same base 3,

set each exponents in the left hand and the right hand sides to be equal.

x=4

Hence, the solution is x=4